Determine the ROR based on 5-year period for a production line if, the costs associated with producing a certain product were as the following:
Initial cost = SR 24,000
AOC = SR 17, 000
Annual revenue = SR 27, 000
Answer #1

Initial cost (P) = 24,000

Net annual benefits (A) = 27000 - 17000 = 10,000

n = 5

P = A(P/A, i, n)

24,000 = 10,000(P/A, i, 5)

(P/A, i, 5) = 24000 / 10000 = 2.4

From the interest table we found that the value of (P/A, i, 5) is at an interest rate of 30% is 2.436 and at 35% it is 2.220.

By interpolation,

i = 0.30 + [(2.436 - 2.4) / (2.436 - 4.220)] * (0.35 - 0.30)

= 0.30 + (0.036 / 0.216) * 0.05

= 0.30 + 0.008

= 0.308 or 30.8%

Thus, the ROR is 30.8%.